• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 06:02:17 on June 8,2010

    Entry number 325210

    keyword=blumi3 HV adjustment

    Blumi3 was close to saturation -- at 230K -- and the 
    Panguin was complaining.
    
    Checked that beam position had not moved.  Is within tolerance.
    
    Reduce Blumi3 HV from 610 to 590.  Signal goes to 177K.